OpenStudy (anonymous):

WILL FAN AND MEDAL! A. Given the equation 2(3x − 4) = 5x + 6, solve for the variable. Explain each step and justify your process. B. Charlie solved a similar equation below. Is Charlie's solution correct? Explain why or why not. 4x − 3 = 2(x − 1) 4x − 3 = 2x + 2 2x − 3 = 2 2x = 6 x =3

OpenStudy (anonymous):

A) 2(3x - 4) = 5x + 6 6x - 8 = 5x + 6 Distributive property. x - 8 = 6 Subtract 5x from both sides. x = 14 Add 8 to both sides. Voila! B) I would probably find the mistake (if there is one) by solving it myself... so I'll do that: 4x - 3 = 2(x + 1) 4x - 3 = 2x + 2 Distributive property (Charlie is right so far) 2x - 3 = 2 Subtract 2x from both sides (Charlie is right so far) 2x = 5 Add three to both sides (Charlie is WRONG here, they multiplied by 3 instead of adding 3) x = 2.5 Divide both sides by 2. So... yeah! Hope I helped, and let me know if you need anything else :D

OpenStudy (anonymous):

2(3x − 4) = 5x + 6 multiply 2 by the (3x -4) 6x -8 = 5x + 6 to get this 6x-5x = 8 + 6 since we have multiplied, get the x together in The left hand side of the equation and constants together on the other side x = 14 you will get x afterwards.

OpenStudy (anonymous):

Thank you so much! @Sionainn

OpenStudy (anonymous):

Anytime, glad I could help :)